# From Entanglement to Alignment: Representation Space Decomposition for Unsupervised Time Series Domain Adaptation

### v1

@@ Conclusions
In this paper, we for the first time shift the paradigm of unsupervised time series domain adaptation from the perspective of explicit representation space decomposition. Building on this insight, we propose DARSD, disentangling domain-invariant patterns from domain-specific artifacts via a learnable invariant subspace basis. Additionally, DARSD decouples the optimization objectives of invariant feature extraction and robust cross-domain feature discriminative aggregation through Adv-LCIB and hybrid contrastive optimization with PPGCE modules. Extensive experiments across 53 scenarios from WISDM, HAR, HHAR, and MFD benchmarks demonstrate its superiority, with DARSD ranking first across all datasets and achieving optimal results in 35 cases. Future work will explore adaptive basis selection to further enhance the DARSD’s applicability across diverse time series domains.
Appendix A Appendix
Appendix B A. Notation
For easier reading, the crucial notations involved in this manuscript are summarized in Table 1.
Table 1: The specific description of crucial notations.
@@ Appendix A Appendix
$\hat{w}\in\mathbb{R}^{m}$
Softmax regularized coordinate of $w^{\textit{inv}}$
$\hat{f}^{s}_{i},\;\hat{f}^{t}_{i}\in\mathbb{R}^{d}$ Reconstructed domain-invariant features
@@
$\sigma_{i}$ confidence score of target feature $\hat{f}^{t}_{i}$
$\hat{f}^{\textit{con}}_{i},\;\hat{f}^{\textit{dis}}_{i}$ Confident and distrusted target feature
@@
$\hat{F}_{t}=\{\hat{f}^{t}_{i}\}_{i=1}^{n_{t}}=\{\hat{F}_{t}^{\textit{con}}\cup\hat{F}_{t}^{\textit{dis}}\}$ Set of reconstructed target features
Appendix C B. Proof of Representation Space Decomposition
@@ B.2. Subspace Representation
$(B^{\textit{inv}})^{\top}B^{\textit{inv}}&=\mathbf{I}_{m},\\ (B^{\textit{spe}})^{\top}B^{\textit{spe}}&=\mathbf{I}_{\bar{m}},\\ (B^{\textit{inv}})^{\top}B^{\textit{spe}}&=\mathbf{0}_{m\times\bar{m}}\end{aligned}$ (3)
@@ B.4. Domain-Invariant Feature Reconstruction
Appendix D C. Detailed Benchmark Description
@@
Figure 1: Extension of Figure 2 in main text. Distributions of $F$ and $\hat{F}$. (a) and (b) illustrate distributions of $F$ and corresponding $\hat{F}$, respectively.
Appendix E D. More Experimental Results
D.1. Hyperparameter Sensitivity Analysis
@@
Figure 2: Hyperparameter sensitivity analysis of $m$ and $\lambda_{2}$ on WISDM 12 $\to$ 19 and HHAR 8 $\to$ 6 scenarios.
@@
Invariant Basis Dimensionality ($m$): Fig. 2a shows that optimal performance occurs at $m=24$ for feature dimensions $d=128$. This indicates that a 24-dimensional subspace provides sufficient capacity to capture essential semantic domain-invariant patterns, enabling effective domain adaptation. When $m<24$, the invariant subspace $B^{\textit{inv}}_{\textit{obs}}$ lacks sufficient capacity to capture semantic complexity, resulting to underfitting. When $m>24$, the enlarged basis becomes harder to train effectively and more susceptible to incorporating domain-specific information, thereby undermining the invariance property.
Adversarial Training Weight ($\lambda_{2}$): Fig. 2b reveals stable performance within $\lambda_{2}\in[0.25,0.50]$, indicating robust optimization dynamics. When $\lambda_{2}<0.25$, insufficient adversarial pressure results in lossy reconstruction of domain-invariant information, causing significant information loss. When $\lambda_{2}>0.50$, excessive adversarial pressure forces the model to incorporate domain-specific information to maintain reconstruction consistency, compromising the invariance property.
@@
D.2. Representation Visualization
Figure 3: t-SNE visualization of DARSD feature distribution in the domain-invariant subspace. Colors distinguish different classes; downward ($\bigtriangledown$) and upward ($\bigtriangleup$) triangles denote source and target features, respectively. Each row shows: (a) source features, (b) target features, and (c) combined source-target feature distribution in the domain-invariant subspace.
Fig. 3 visualizes distributions of source and target features in the invariant subspace using t-SNE, providing compelling evidence for DARSD’s effectiveness. Source features (Fig. 3a) form well-separated semantic clusters with distinct inter-class boundaries. Target features (Fig. 3b) display a remarkably similar geometric structure, where cluster shapes and relative positions mirror those of the source distribution. This structural preservation indicates that our Adv-LCIB module successfully extracts transferable domain-invariant semantic patterns while filtering domain-specific artifacts. Fig. 3c displays the substantial spatial overlap between source and target features of identical classes, alongside maintained inter-class separation. This confirms that our DARSD framework achieves dual objectives of domain invariance and semantic discriminability. Notably, explainable feature visualization validates that learned representations are truly domain-agnostic, supporting our theoretical representation space decomposition framework and the effectiveness of PPGCE in the domain-invariant subspace.
D.3. Convergence Efficiency
Figure 4: Comparisons of convergence efficiency of DARSD and CLUDA. The red and blue arrows indicate convergence trends of DARSD and CLUDA, respectively.
Fig. 4 compares the convergence efficiency of DARSD and CLUDA across representative scenarios. DARSD consistently exhibits faster convergence, typically reaching stable performance within 50-80 epochs compared to CLUDA’s 100-150 epochs.
@@ D.3. Convergence Efficiency
Appendix F E. Related Works
@@ E.1. Unsupervised Domain Adaptation
Unsupervised domain adaptation (UDA) seeks to transfer knowledge from an annotated source domain to an unlabeled target domain, where two domains have related yet different distributions. Despite significant progress, existing approaches suffer from a fundamental limitation: they treat representations as monolithic entities and strive to align entire feature distributions, failing to recognize that only part of components contain transferable knowledge. They can be categorized into three groups: (1) Adversarial Training Methods represent the earliest paradigm for UDA (Tzeng et al. 2017), employing domain discriminators to learn domain-confusing representations (Ozyurt, Feuerriegel, and Zhang 2023). VRADA (Purushotham et al. 2017) and CDAN (Long et al. 2018) utilize adversarial training to align marginal distributions, while CoDATS (Wilson, Doppa, and Cook 2020) extends this framework to time series with multi-source adaptation. However, these methods suffer from a critical flaw: the adversarial game fails to distinguish between transferable semantic content and domain-specific noise, resulting in the elimination of essential information required for accurate classification. (2) Metric Learning Approaches aim to mitigate domain shift by explicitly minimizing statistical distances between domain distributions. Methods such as DDC (Tzeng et al. 2014) and DeepCORAL (Sun and Saenko 2016) align second-order statistics, while more advanced techniques like HoMM (Chen et al. 2020) perform high-order moment matching. Despite their theoretical grounding, these methods rely on an untenable assumption: that entire feature vectors should be aligned. This overlooks the compositional nature of representations, where only specific dimensions encode transferable knowledge. The introduction of domain-specific artifacts poisons the statistical properties of high-dimensional features, distorting their transferability across domains. (3) Self-supervised Methods leverage contrastive learning guided by pretext tasks to acquire domain-invariant representations. Approaches like CLUDA (Ozyurt, Feuerriegel, and Zhang 2023) progressively extracts features via a MoCo architecture enhanced with adversarial training, while other studies (Xu et al. 2025; He et al. 2023; Jin et al. 2022) employ diverse pretext tasks to promote cross-domain consistency. Critically, the design of pretext tasks directly governs the quality and transferability of the resulting domain-invariant information.
@@
Despite their methodological differences, all existing approaches share a common philosophical limitation: they rely on implicit alignment mechanisms. Whether via adversarial confusion, statistical distance minimization, or contrastive consistency, these methods presume that global feature alignment will simultaneously achieve domain invariant and preserve semantic content. This lack of explicit control over information preservation constitutes a critical weakness: adaptation occurs without principled mechanisms to discern what knowledge is retained versus discarded during feature extraction.

### v2

@@ Conclusions
In this paper, we for the first time shift the paradigm of unsupervised time series domain adaptation from the perspective of explicit representation space decomposition. Building on this insight, we propose DARSD, disentangling domain-invariant patterns from domain-specific artifacts via a learnable invariant subspace basis. Additionally, DARSD decouples the optimization objectives of invariant feature extraction and robust cross-domain feature discriminative aggregation through Adv-LCIB and hybrid contrastive optimization with PPGCE modules. Extensive experiments conducted on four benchmarks demonstrate its superiority. Comprehensive studies prove the rationality of individual elements in DARSD and their synergy to achieve optimal performance. Future work will explore adaptive basis selection to further enhance the DARSD’s applicability across diverse time series domains.
Appendix
Appendix A A. Notation
For easier reading, the crucial notations involved in this manuscript are summarized in Appendix Table 1.
@@
$\hat{w}\in\mathbb{R}^{m}$ Softmax regularized coordinate
$\hat{f}^{s}_{i},\;\hat{f}^{t}_{i}\in\mathbb{R}^{d}$
Reconstructed domain-invariant
features
@@
$\sigma_{i}$ Confidence score of $\hat{f}^{t}_{i}$
$\hat{f}^{\textit{con}}_{i}$ Confident target feature
$\hat{f}^{\textit{dis}}_{i}$ Distrusted target feature
@@
$\hat{F}_{t}=\{\hat{F}_{t}^{\textit{con}}\cup\hat{F}_{t}^{\textit{dis}}\}$ Set of reconstructed target features
Appendix Table 1: Specific description of crucial notations.
Appendix Figure 1: Extension of Figure 3 in main text. Distributions of $F$ and $\hat{F}$. (a) and (b) illustrate distributions of $F$ and corresponding $\hat{F}$, respectively.
Appendix B B. Proof of Representation Space Decomposition
@@ B.2. Subspace Representation
$\begin{aligned} (B^{\textit{inv}})^{\top}B^{\textit{inv}}&=\mathbf{I}_{m},\\ (B^{\textit{spe}})^{\top}B^{\textit{spe}}&=\mathbf{I}_{\bar{m}},\\ (B^{\textit{inv}})^{\top}B^{\textit{spe}}&=\mathbf{0}_{m\times\bar{m}}\end{aligned}$ (3)
@@ B.4. Domain-Invariant Feature Reconstruction
Appendix C C. Detailed Benchmark Description
@@
Appendix Figure 2: Hyperparameter sensitivity analysis of $m$ and $\lambda_{2}$ on WISDM 12 $\to$ 19 and HHAR 8 $\to$ 6 scenarios.
Appendix D D. More Experimental Results
D.1. Macro-F1 Results
Src $\to$ Trg DDC D-CORAL VRADA CDAN CAN CoDATS HoMM MMDA AdvSKM DSAN CLUDA CADT DARSD
WISDM; Out of 19 scenarios, DARSD achieves the 15 optimal results and 3 runner-up positions.
10 $\to$ 14 0.541 0.337 0.591 0.266 0.608 0.477 0.667 0.469 0.471 0.337 0.510 0.599 0.691
10 $\to$ 21 0.694 0.250 0.125 0.118 0.250 0.390 0.313 0.390 0.694 0.106 0.250 0.518 0.392
12 $\to$ 7 0.632 0.486 0.437 0.546 0.636 0.612 0.442 0.539 0.655 0.574 0.678 0.620 0.834
12 $\to$ 19 0.396 0.317 0.410 0.298 0.508 0.456 0.281 0.233 0.510 0.518 0.532 0.518 0.883
15 $\to$ 29 0.299 0.265 0.391 0.238 0.372 0.136 0.208 0.133 0.299 0.182 0.482 0.441 0.529
18 $\to$ 20 0.383 0.379 0.578 0.600 0.389 0.427 0.421 0.280 0.348 0.268 0.673 0.702 0.658
19 $\to$ 2 0.459 0.501 0.615 0.312 0.327 0.403 0.522 0.306 0.460 0.428 0.458 0.555 0.664
21 $\to$ 16 0.537 0.669 0.542 0.404 0.641 0.410 0.669 0.310 0.523 0.568 0.595 0.850 0.787
25 $\to$ 22 0.273 0.241 0.506 0.401 0.241 0.200 0.307 0.159 0.309 0.341 0.375 0.500 0.610
26 $\to$ 2 0.414 0.618 0.517 0.404 0.362 0.598 0.519 0.453 0.463 0.424 0.701 0.610 0.731
28 $\to$ 2 0.484 0.495 0.473 0.400 0.412 0.492 0.511 0.430 0.484 0.451 0.710 0.606 0.656
28 $\to$ 20 0.571 0.620 0.672 0.605 0.655 0.578 0.699 0.537 0.557 0.615 0.703 0.726 0.728
HAR; Out of 11 scenarios, DARSD achieves the 6 optimal results and 4 runner-up positions.
18 $\to$ 21 0.439 0.760 0.697 0.450 0.449 0.857 0.803 0.380 0.439 0.588 0.920 0.929 0.948
20 $\to$ 6 0.643 0.776 0.619 0.407 0.593 0.659 0.935 0.407 0.586 0.588 0.959 0.782 0.967
23 $\to$ 13 0.398 0.586 0.683 0.298 0.509 0.439 0.539 0.518 0.401 0.723 0.720 0.743 0.764
24 $\to$ 12 0.724 0.765 0.658 0.758 0.872 0.709 0.893 0.448 0.736 0.826 0.866 0.768 0.887
Avg. Rank 10.2 6.55 5.55 11.0 8.45 8.18 4.82 12.0 8.36 6.36 3.73 3.64 1.64
HHAR; Out of 13 scenarios, DARSD achieves the 9 optimal results and 4 runner-up positions.
5 $\to$ 2 0.311 0.330 0.327 0.324 0.338 0.209 0.335 0.278 0.229 0.232 0.299 0.318 0.340
7 $\to$ 0 0.194 0.198 0.169 0.454 0.178 0.282 0.232 0.095 0.189 0.247 0.438 0.304 0.527
7 $\to$ 2 0.233 0.295 0.525 0.333 0.250 0.217 0.258 0.235 0.203 0.179 0.332 0.390 0.392
7 $\to$ 4 0.577 0.608 0.658 0.344 0.736 0.489 0.750 0.539 0.543 0.522 0.604 0.850 0.813
7 $\to$ 5 0.488 0.496 0.529 0.480 0.546 0.374 0.323 0.283 0.487 0.649 0.626 0.645 0.680
8 $\to$ 6 0.568 0.728 0.723 0.334 0.765 0.701 0.745 0.514 0.734 0.467 0.837 0.773 0.844
Avg. Rank 9.38 7.31 6.85 6.77 7.54 9.00 6.62 10.7 9.15 9.00 3.38 3.85 1.31
MFD; Out of 10 scenarios, DARSD achieves the 5 optimal results and 2 runner-up positions.
1 $\to$ 0 0.500 0.514 0.427 0.555 0.507 0.614 0.506 0.515 0.500 0.484 0.494 0.709 0.663
1 $\to$ 2 0.749 0.704 0.500 0.881 0.572 0.919 0.747 0.773 0.734 0.732 0.574 0.680 0.776
0 $\to$ 1 0.524 0.504 0.455 0.599 0.518 0.607 0.509 0.593 0.468 0.374 0.510 0.515 0.722
Avg. Rank 5.70 8.90 12.4 3.10 10.8 3.10 7.80 4.80 9.10 8.50 7.50 6.90 2.10
Appendix Table 2: Extension of Table 2 in main text. Macro-F1 results on WISDM, HAR, HHAR, and MFD benchmarks. Each result is the mean over ten random initializations. Src and Trg denote the source and target subsets. The best two results are highlighted by 1st and 2nd.
The Appendix Table 2 displays the remaining part of Table 2 in main text.
D.2. Hyperparameter Sensitivity Analysis
@@
Invariant Basis Dimensionality ($m$): Appendix Fig. 2a shows that optimal performance occurs at $m=24$ for feature dimensions $d=128$. This indicates that a 24-dimensional subspace provides sufficient capacity to capture essential semantic domain-invariant patterns, enabling effective domain adaptation. When $m<24$, the invariant subspace $B^{\textit{inv}}_{\textit{obs}}$ lacks sufficient capacity to capture semantic complexity, resulting in underfitting. When $m>24$, the enlarged basis becomes harder to train effectively and more susceptible to incorporating domain-specific information, thereby undermining the invariance property.
Adversarial Training Weight ($\lambda_{2}$): Appendix Fig. 2b reveals stable performance within $\lambda_{2}\in[0.25,0.50]$, indicating robust optimization dynamics. When $\lambda_{2}<0.25$, insufficient adversarial pressure results in lossy reconstruction of domain-invariant information, causing significant information loss. When $\lambda_{2}>0.50$, excessive adversarial pressure forces the model to incorporate domain-specific information to maintain reconstruction consistency, compromising the invariance property.
@@
D.3. Representation Visualization
Appendix Figure 3: t-SNE visualization of DARSD feature distribution in the domain-invariant subspace. Colors distinguish different classes; downward ($\bigtriangledown$) and upward ($\bigtriangleup$) triangles denote source and target features, respectively. Each row shows: (a) source features, (b) target features, and (c) combined source-target feature distribution in the domain-invariant subspace.
Appendix Fig. 3 visualizes distributions of source and target features in the invariant subspace using t-SNE, providing compelling evidence for DARSD’s effectiveness. Source features (Appendix Fig. 3a) form well-separated semantic clusters with distinct inter-class boundaries. Target features (Appendix Fig. 3b) display a remarkably similar geometric structure, where cluster shapes and relative positions mirror those of the source distribution. This structural preservation indicates that our Adv-LCIB module successfully extracts transferable domain-invariant semantic patterns while filtering domain-specific artifacts. Appendix Fig. 3c displays the substantial spatial overlap between source and target features of identical classes, alongside maintained inter-class separation. This confirms that our DARSD framework achieves dual objectives of domain invariance and semantic discriminability. Notably, explainable feature visualization validates that learned representations are truly domain-agnostic, supporting our theoretical representation space decomposition framework and the effectiveness of PPGCE in the domain-invariant subspace.
D.4. Convergence Efficiency
Appendix Figure 4: Comparisons of convergence efficiency of DARSD and CLUDA. The red and blue arrows indicate convergence trends of DARSD and CLUDA, respectively.
Appendix Fig. 4 compares the convergence efficiency of DARSD and CLUDA across representative scenarios. DARSD consistently exhibits faster convergence, typically reaching stable performance within 50-80 epochs compared to CLUDA’s 100-150 epochs.
@@ D.4. Convergence Efficiency
Appendix E E. Related Works
@@ E.1. Unsupervised Domain Adaptation
Unsupervised domain adaptation (UDA) seeks to transfer knowledge from an annotated source domain to an unlabeled target domain, where two domains have related yet different distributions. Despite significant progress, existing approaches suffer from a fundamental limitation: they treat representations as monolithic entities and strive to align entire feature distributions, failing to recognize that only part of components contain transferable knowledge. They can be categorized into three groups: (1) Adversarial Training Methods represent the earliest paradigm for UDA (Tzeng et al. 2017), employing domain discriminators to learn domain-confusing representations. MBAN (Wang, Yang, and Yu 2024) and CPDIC (Huang et al. 2024) utilize adversarial training to align marginal distributions, while CoDATS (Wilson, Doppa, and Cook 2020) extends this framework to time series with multi-source adaptation. However, these methods suffer from a critical flaw: the adversarial game fails to distinguish between transferable semantic content and domain-specific noise, resulting in the elimination of essential information required for accurate classification. (2) Metric Learning Approaches aim to mitigate domain shift by explicitly minimizing statistical distances between domain distributions. Methods proposed in  (Zha et al. 2025; Zhu et al. 2021) align second-order statistics in Hilbert space, while other techniques like HoMM (Chen et al. 2020) perform high-order moment matching. Despite their theoretical grounding, these methods rely on an untenable assumption: that entire feature vectors should be aligned. This overlooks the compositional nature of representations, where only specific dimensions encode transferable knowledge. The introduction of domain-specific artifacts poisons the statistical properties of high-dimensional features, distorting their transferability across domains. (3) Self-supervised Methods leverage contrastive learning guided by pretext tasks to acquire domain-invariant representations. Approaches like CLUDA (Ozyurt, Feuerriegel, and Zhang 2023) progressively extract features via a MoCo architecture enhanced with adversarial training, while other studies (Xu et al. 2025; He et al. 2023; Jin et al. 2022) employ diverse pretext tasks to promote cross-domain consistency. Critically, the design of pretext tasks directly governs the quality and transferability of the resulting domain-invariant information.
@@
Despite their methodological differences, all existing approaches share a common philosophical limitation: they rely on implicit alignment mechanisms. Whether via adversarial confusion, statistical distance minimization, or contrastive consistency, these methods presume that global feature alignment will simultaneously achieve domain invariance and preserve semantic content. This lack of explicit control over information preservation constitutes a critical weakness: adaptation occurs without principled mechanisms to discern what knowledge is retained versus discarded during feature extraction.